OpenStudy (anonymous):

The median size of a new single family home grew from 2063ft^2 in 1996 to 2523ft^2 in 2007. What is the percent of increase?

OpenStudy (lgbasallote):

2523 - 2063 = 460 460/2063 x 100 = % increase 22.30 % is the percent increase

OpenStudy (anonymous):

The number on two consecutively numbered gym lockers have a sum of159. what are the locker numbers. Use comma to separate the answer

OpenStudy (lgbasallote):

let x = first locker x+1 = second locker x + x + 1 = 159 2x + 1 = 159 2x = 158 x = 79 x+1 = 80

OpenStudy (anonymous):

it should be two numbers

OpenStudy (lgbasallote):

yeah first locker is 79...second locker is 80

OpenStudy (anonymous):

Sarah investment in stock grew 22% to $366. how much did she invest?